# A Framework for Objective-Driven Dynamical Stochastic Fields

### v1

@@ Appendix A Proofs in Section 2
$\langle{\hat{\bm{e}}}_{\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}},{{\mathbf{G}}}{\hat{\bm{e}}}_{\alpha\beta\mu\nu}\rangle\quad\text{and}\quad\langle{\hat{\bm{e}}}_{\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}},({{\mathbf{M}}}+{{\mathbf{N}}}){\hat{\bm{e}}}_{\alpha\beta\mu\nu}\rangle$ (A.19)
@@
$\langle{\hat{\bm{e}}}_{\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}},{{\mathbf{G}}}{\hat{\bm{e}}}_{\alpha\beta\mu\nu}\rangle\cdot{\Delta t}$ $=p_{{\Delta t}}(\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}\mid\alpha\beta\mu\nu)-\delta_{\alpha\beta\mu\nu}^{\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}}+o({\Delta t})$ (A.20)
@@
$=(\langle{\hat{\bm{e}}}_{\alpha^{\prime}\mu^{\prime}},{{\mathbf{M}}}{\hat{\bm{e}}}_{\alpha\beta\mu}\rangle\cdot{\Delta t}+\delta_{\alpha\mu}^{\alpha^{\prime}\mu^{\prime}})\cdot(\langle{\hat{\bm{e}}}_{\beta^{\prime}\nu^{\prime}},{{\mathbf{N}}}{\hat{\bm{e}}}_{\alpha\beta\nu}\rangle\cdot{\Delta t}+\delta_{\beta\nu}^{\beta^{\prime}\nu^{\prime}})-\delta_{\alpha\beta\mu\nu}^{\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}}+o({\Delta t})$ (A.23)
$=\langle{\hat{\bm{e}}}_{\alpha^{\prime}\mu^{\prime}},{{\mathbf{M}}}{\hat{\bm{e}}}_{\alpha\beta\mu}\rangle\cdot{\Delta t}\cdot\delta_{\beta\nu}^{\beta^{\prime}\nu^{\prime}}+\langle{\hat{\bm{e}}}_{\beta^{\prime}\nu^{\prime}},{{\mathbf{N}}}{\hat{\bm{e}}}_{\alpha\beta\nu}\rangle\cdot{\Delta t}\cdot\delta_{\alpha\mu}^{\alpha^{\prime}\mu^{\prime}}+o({\Delta t}).$ (A.24)
@@
It left to convert the above to the expression with their embeddings ${{\mathbf{M}}},{{\mathbf{N}}}$. By definition of the embeddings (Definition 2.6) we obtain
@@
$=\langle{\hat{\bm{e}}}_{\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}},({{\mathbf{M}}}{\hat{\bm{e}}}_{\alpha\beta\mu})\otimes{\hat{\bm{e}}}_{\beta\nu}\rangle\cdot{\Delta t}+\langle{\hat{\bm{e}}}_{\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}},({{\mathbf{N}}}{\hat{\bm{e}}}_{\alpha\beta\nu})\otimes{\hat{\bm{e}}}_{\alpha\mu}\rangle\cdot{\Delta t}+o({\Delta t})$ (A.26)
$=\langle{\hat{\bm{e}}}_{\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}},({{\mathbf{M}}}+{{\mathbf{N}}}){\hat{\bm{e}}}_{\alpha\beta\mu\nu}\rangle\cdot{\Delta t}+o({\Delta t}).$ (A.27)
@@
$\langle{\hat{\bm{e}}}_{\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}},{{\mathbf{G}}}{\hat{\bm{e}}}_{\alpha\beta\mu\nu}\rangle\cdot{\Delta t}=\langle{\hat{\bm{e}}}_{\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}},({{\mathbf{M}}}+{{\mathbf{N}}}){\hat{\bm{e}}}_{\alpha\beta\mu\nu}\rangle\cdot{\Delta t}+o({\Delta t}).$ (A.28)
@@
$\langle{\hat{\bm{e}}}_{\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}},{{\mathbf{G}}}{\hat{\bm{e}}}_{\alpha\beta\mu\nu}\rangle=\langle{\hat{\bm{e}}}_{\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}},({{\mathbf{M}}}+{{\mathbf{N}}}){\hat{\bm{e}}}_{\alpha\beta\mu\nu}\rangle.$ (A.29)
@@
$\langle{\hat{\bm{e}}}_{{\omega}^{\prime}},{{\mathbf{G}}}{\hat{\bm{e}}}_{{\omega}}\rangle\cdot{\Delta t}$ $=p_{{\Delta t}}({\omega}^{\prime}\mid{\omega})-\delta_{{\omega}}^{{\omega}^{\prime}}+o({\Delta t}),$ (A.32)
$\langle{\hat{\bm{e}}}_{{\omega}^{\prime}_{|{\Omega}_{x}}},{{\mathbf{G}}}(x){\hat{\bm{e}}}_{{\omega}_{|{{\mathcal{U}}}_{x}}}\rangle\cdot{\Delta t}$ $=p_{{\Delta t}}({\omega}^{\prime}_{|{\Omega}_{x}}\mid{\omega}_{|{{\mathcal{U}}}_{x}})-\delta_{{\omega}_{|{\Omega}_{x}}}^{{\omega}^{\prime}_{|{\Omega}_{x}}}+o({\Delta t}).$ (A.33)
@@
$\langle{\hat{\bm{e}}}_{{\omega}^{\prime}},{{\mathbf{G}}}{\hat{\bm{e}}}_{{\omega}}\rangle\quad\text{and}\quad\langle{\hat{\bm{e}}}_{{\omega}^{\prime}},(\sum_{x}{{\mathbf{G}}}(x)){\hat{\bm{e}}}_{{\omega}}\rangle$ (A.35)
@@
$\langle{\hat{\bm{e}}}_{{\omega}^{\prime}},{{\mathbf{G}}}{\hat{\bm{e}}}_{{\omega}}\rangle\cdot{\Delta t}$ $=p_{{\Delta t}}({\omega}^{\prime}\mid{\omega})-\delta_{{\omega}}^{{\omega}^{\prime}}+o({\Delta t})=\prod_{x}p_{\Delta t}({\omega}^{\prime}_{|{\Omega}_{x}}\mid{\omega}_{|{{\mathcal{U}}}_{x}})-\delta_{{\omega}}^{{\omega}^{\prime}}+o({\Delta t}).$ (A.36)
@@
$=\prod_{x}\left(\langle{\hat{\bm{e}}}_{{\omega}^{\prime}_{|{\Omega}_{x}}},{{\mathbf{G}}}(x){\hat{\bm{e}}}_{{\omega}_{|{{\mathcal{U}}}_{x}}}\rangle\cdot{\Delta t}+\delta_{{\omega}_{|{\Omega}_{x}}}^{{\omega}^{\prime}_{|{\Omega}_{x}}}\right)-\delta_{{\omega}}^{{\omega}^{\prime}}+o({\Delta t})$ (A.38)
$=\sum_{x}\langle{\hat{\bm{e}}}_{{\omega}^{\prime}_{|{\Omega}_{x}}},{{\mathbf{G}}}(x){\hat{\bm{e}}}_{{\omega}_{|{{\mathcal{U}}}_{x}}}\rangle\cdot{\Delta t}\cdot\delta_{{\omega}_{|{\Omega}/{\Omega}_{x}}}^{{\omega}^{\prime}_{|{\Omega}/{\Omega}_{x}}}+o({\Delta t}).$ (A.39)
@@
$=\sum_{x}\langle{\hat{\bm{e}}}_{{\omega}^{\prime}_{|{\Omega}_{x}}}\otimes{\hat{\bm{e}}}_{{\omega}^{\prime}_{|{\Omega}/{\Omega}_{x}}},({{\mathbf{G}}}(x){\hat{\bm{e}}}_{{\omega}_{|{{\mathcal{U}}}_{x}}})\otimes{\hat{\bm{e}}}_{{\omega}_{|{\Omega}/{\Omega}_{x}}}\rangle\cdot{\Delta t}+o({\Delta t})$ (A.41)
$=\sum_{x}\langle{\hat{\bm{e}}}_{{\omega}^{\prime}},{{\mathbf{G}}}(x){\hat{\bm{e}}}_{{\omega}}\rangle\cdot{\Delta t}+o({\Delta t}).$ (A.42)
@@
$\langle{\hat{\bm{e}}}_{{\omega}^{\prime}},{{\mathbf{G}}}{\hat{\bm{e}}}_{{\omega}}\rangle\cdot{\Delta t}=\langle{\hat{\bm{e}}}_{{\omega}^{\prime}},(\sum_{x}{{\mathbf{G}}}(x)){\hat{\bm{e}}}_{{\omega}_{|{{\mathcal{U}}}_{x}}}\rangle\cdot{\Delta t}+o({\Delta t}).$ (A.43)
@@
$\langle{\hat{\bm{e}}}_{{\omega}^{\prime}},{{\mathbf{G}}}{\hat{\bm{e}}}_{{\omega}}\rangle=\langle{\hat{\bm{e}}}_{{\omega}^{\prime}},(\sum_{x}G(x)){\hat{\bm{e}}}_{{\omega}}\rangle.$ (A.44)
@@
where the commutator $[{{\mathbf{G}}}(x),{{\mathbf{G}}}(x^{\prime})]={{\mathbf{G}}}(x){{\mathbf{G}}}(x^{\prime})-{{\mathbf{G}}}(x^{\prime}){{\mathbf{G}}}(x)$.
@@
${{\mathbf{G}}}(x^{\prime}){{\mathbf{G}}}(x){\hat{\bm{e}}}_{{\omega}}$ $={{\mathbf{G}}}(x^{\prime})\left(({{\mathbf{G}}}(x){\hat{\bm{e}}}_{{\omega}_{|{{\mathcal{U}}}_{x}}})\otimes{\hat{\bm{e}}}_{{\omega}_{|{\Omega}/{\Omega}_{x}}}\right)$ (A.47)
$=({{\mathbf{G}}}(x){\hat{\bm{e}}}_{{\omega}_{|{{\mathcal{U}}}_{x}}})\otimes({{\mathbf{G}}}(x^{\prime}){\hat{\bm{e}}}_{{\omega}_{|{{\mathcal{U}}}_{x^{\prime}}}})\otimes{\hat{\bm{e}}}_{{\omega}_{|{\Omega}/({\Omega}_{x}\times{\Omega}_{x^{\prime}})}}$ (A.48)
$={{\mathbf{G}}}(x){{\mathbf{G}}}(x^{\prime}){\hat{\bm{e}}}_{{\omega}}.$ (A.49)
@@
$L(w_{t},w_{t}^{+})=-\left(\sum_{x}G(x)^{{\omega}_{t}}_{{\omega}_{t}}\right)-\delta({{\omega}_{t}^{+}\neq{\omega}_{t}})\cdot\log\left(\sum_{x}G(x)^{{\omega}_{t}^{+}}_{{\omega}_{t}}\right).$ (A.51)
@@
$\delta({{\omega}_{t}^{+}\neq{\omega}_{t}})\cdot\log\left(\sum_{x}G(x)^{{\omega}_{t}^{+}}_{{\omega}_{t}}\right)$ $=\sum_{x}\delta({{\omega}_{t}^{+}}_{|{\Omega}_{x}}\neq{{\omega}_{t}}_{|{\Omega}_{x}})\cdot\log\left(G(x)^{{{\omega}_{t}^{+}}_{|{\Omega}_{x}}}_{{{\omega}_{t}}_{|{{\mathcal{U}}}_{x}}}\right).$ (A.52)
@@
$p_{T}({\omega}^{\prime}\mid{\omega})=\langle{\hat{\bm{e}}}_{{\omega}^{\prime}},e^{{{\mathbf{G}}}T}{\hat{\bm{e}}}_{{\omega}}\rangle=\int{\mathcal{D}}{\omega}\ \exp\left\{-\sum_{x}\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ L(w_{t,x},w_{t,x}^{+})\right\},$ (A.54)
@@
$p_{t}({\omega}^{\prime}\mid{\omega})$ $=\sum_{k=0}^{N}\sum_{{\omega}_{0:T,{\Delta t},k}}^{{\omega}\to{\omega}^{\prime}}\prod_{n=1}^{N}\langle{\hat{\bm{e}}}_{{\omega}_{t_{n}}},e^{{{\mathbf{G}}}{\Delta t}}{\hat{\bm{e}}}_{{\omega}_{t_{n-1}}}\rangle$ (A.55)
$=\sum_{k=0}^{N}\sum_{{\omega}_{0:T,{\Delta t},k}}^{{\omega}\to{\omega}^{\prime}}\cdot({\Delta t})^{k}\cdot\frac{1}{({\Delta t})^{k}}\prod_{n=1}^{N}\langle{\hat{\bm{e}}}_{{\omega}_{t_{n}}},e^{{{\mathbf{G}}}{\Delta t}}{\hat{\bm{e}}}_{{\omega}_{t_{n-1}}}\rangle$ (A.56)
$=\sum_{k=0}^{N}\sum_{{\omega}_{0:T,{\Delta t},k}}^{{\omega}\to{\omega}^{\prime}}\cdot({\Delta t})^{k}\cdot\exp\left\{-k\log\left({\Delta t}\right)+\sum_{n=1}^{N}\log\ \langle{\hat{\bm{e}}}_{{\omega}_{t_{n}}},e^{{{\mathbf{G}}}{\Delta t}}{\hat{\bm{e}}}_{{\omega}_{t_{n-1}}}\rangle\right\}$ (A.57)
@@
$-k\log\left({\Delta t}\right)+\sum_{n=1}^{N}\log\ \langle{\hat{\bm{e}}}_{{\omega}_{t_{n}}},e^{{{\mathbf{G}}}{\Delta t}}{\hat{\bm{e}}}_{{\omega}_{t_{n-1}}}\rangle,$ (A.58)
@@
$e^{{{\mathbf{G}}}{\Delta t}}=1+{{\mathbf{G}}}{\Delta t}+o({\Delta t})\quad\text{and}\quad{{\mathbf{G}}}={\hat{\bm{e}}}_{{\omega}^{\prime}}G^{{\omega}^{\prime}}_{{\omega}}{{\check{\bm{\eta}}}}^{{\omega}}.$ (A.59)
@@
$\log\ \langle{\hat{\bm{e}}}_{{\omega}_{t_{n}}},e^{{{\mathbf{G}}}{\Delta t}}{\hat{\bm{e}}}_{{\omega}_{t_{n-1}}}\rangle=\log\left(1+G^{{\omega}_{t_{n}}}_{{\omega}_{t_{n-1}}}\cdot{\Delta t}+o({\Delta t})\right)=G^{{\omega}_{t_{n}}}_{{\omega}_{t_{n-1}}}\cdot{\Delta t}+o({\Delta t}).$ (A.60)
For the second case, i.e., ${\omega}_{t_{n}}\neq{\omega}_{t_{n-1}}$, there are exactly $k$ jumps for the given path. Therefore, we distribute the $-k\log({\Delta t})$ into this case and obtain
$-\log({\Delta t})+\log\ \langle{\hat{\bm{e}}}_{{\omega}_{t_{n}}},e^{{{\mathbf{G}}}{\Delta t}}{\hat{\bm{e}}}_{{\omega}_{t_{n-1}}}\rangle$ $=\log\left(\langle{\hat{\bm{e}}}_{{\omega}_{t_{n}}},e^{{{\mathbf{G}}}{\Delta t}}{\hat{\bm{e}}}_{{\omega}_{t_{n-1}}}\rangle/{\Delta t}\right)$ (A.61)
@@
When taking ${\Delta t}\to 0$, or equivalently $N=T/{\Delta t}\to\infty$, the $\sum_{n=1}^{N}{\Delta t}$ above becomes an integral $\int_{0}^{T}{\mathop{}\!\mathrm{d}t}$. The last term becomes $o({\Delta t})/{\Delta t}\to 0$. Then, ${\omega}_{t}$ is discontinuous w.r.t. time but with only finite discontinuous point, i.e., jumps between configurations. Since there is only finite jumps, the first term becomes just $G^{{\omega}_{t}}_{{\omega}_{t}}$, i.e.,
@@
$\lim_{{\Delta t}\to 0}\sum_{n=1}^{N}{\Delta t}\cdot(1-\delta_{{\omega}_{t_{n-1}}}^{{\omega}_{t_{n}}})\log\left(G^{{\omega}_{t_{n}}}_{{\omega}_{t_{n-1}}}\right)/{\Delta t}=\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ \delta({{\omega}_{t}^{+}\neq{\omega}_{t}})\cdot\log\left(G^{{\omega}_{t}^{+}}_{{\omega}_{t}}\right).$ (A.66)
@@
$\lim_{{\Delta t}\to 0}-k\log\left({\Delta t}\right)+\sum_{n=1}^{N}\log\ \langle{\hat{\bm{e}}}_{{\omega}_{t_{n}}},e^{{{\mathbf{G}}}{\Delta t}}{\hat{\bm{e}}}_{{\omega}_{t_{n-1}}}\rangle=-\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ L(w_{t},w_{t}^{+}).$ (A.67)
Thus, adopting the path integral notations, we find the path integral formulation of the propagation probability
$p_{T}({\omega}^{\prime}\mid{\omega})$ $=\lim_{{\Delta t}\to 0}\sum_{k=0}^{N}\sum_{{\omega}_{0:T,{\Delta t},k}}^{{\omega}\to{\omega}^{\prime}}\cdot({\Delta t})^{k}\cdot\exp\left\{-k\log\left({\Delta t}\right)+\sum_{n=1}^{N}\log\ \langle{\hat{\bm{e}}}_{{\omega}_{t_{n}}},e^{{{\mathbf{G}}}{\Delta t}}{\hat{\bm{e}}}_{{\omega}_{t_{n-1}}}\rangle\right\}$ (A.68)

### v2

@@ Appendix A Proofs in Section 2
$\langle\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}\rvert{{\mathbf{G}}}\lvert\alpha\beta\mu\nu\rangle\quad\text{and}\quad\langle\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}\rvert({{\mathbf{M}}}+{{\mathbf{N}}})\lvert\alpha\beta\mu\nu\rangle$ (A.19)
@@
$\langle\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}\rvert{{\mathbf{G}}}\lvert\alpha\beta\mu\nu\rangle\cdot{\Delta t}$ $=p_{{\Delta t}}(\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}\mid\alpha\beta\mu\nu)-\delta_{\alpha\beta\mu\nu}^{\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}}+o({\Delta t})$ (A.20)
@@
$=(\langle\alpha^{\prime}\mu^{\prime}\rvert{{\mathbf{M}}}\lvert\alpha\beta\mu\rangle\cdot{\Delta t}+\delta_{\alpha\mu}^{\alpha^{\prime}\mu^{\prime}})\cdot(\langle\beta^{\prime}\nu^{\prime}\rvert{{\mathbf{N}}}\lvert\alpha\beta\nu\rangle\cdot{\Delta t}+\delta_{\beta\nu}^{\beta^{\prime}\nu^{\prime}})-\delta_{\alpha\beta\mu\nu}^{\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}}+o({\Delta t})$ (A.23)
$=\langle\alpha^{\prime}\mu^{\prime}\rvert{{\mathbf{M}}}\lvert\alpha\beta\mu\rangle\cdot{\Delta t}\cdot\delta_{\beta\nu}^{\beta^{\prime}\nu^{\prime}}+\langle\beta^{\prime}\nu^{\prime}\rvert{{\mathbf{N}}}\lvert\alpha\beta\nu\rangle\cdot{\Delta t}\cdot\delta_{\alpha\mu}^{\alpha^{\prime}\mu^{\prime}}+o({\Delta t}).$ (A.24)
@@
It remains to convert the above to the expression with their embeddings ${{\mathbf{M}}},{{\mathbf{N}}}$. By definition of the embeddings (Definition 2.6) we obtain
@@
$=\langle\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}\rvert(({{\mathbf{M}}}\lvert\alpha\beta\mu\rangle)\otimes\lvert\beta\nu\rangle)\cdot{\Delta t}+\langle\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}\rvert(({{\mathbf{N}}}\lvert\alpha\beta\nu\rangle)\otimes\lvert\alpha\mu\rangle)\cdot{\Delta t}+o({\Delta t})$ (A.26)
$=\langle\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}\rvert({{\mathbf{M}}}+{{\mathbf{N}}})\lvert\alpha\beta\mu\nu\rangle\cdot{\Delta t}+o({\Delta t}).$ (A.27)
@@
$\langle\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}\rvert{{\mathbf{G}}}\lvert\alpha\beta\mu\nu\rangle\cdot{\Delta t}=\langle\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}\rvert({{\mathbf{M}}}+{{\mathbf{N}}})\lvert\alpha\beta\mu\nu\rangle\cdot{\Delta t}+o({\Delta t}).$ (A.28)
@@
$\langle\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}\rvert{{\mathbf{G}}}\lvert\alpha\beta\mu\nu\rangle=\langle\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}\rvert({{\mathbf{M}}}+{{\mathbf{N}}})\lvert\alpha\beta\mu\nu\rangle.$ (A.29)
@@
$\langle{\omega}^{\prime}\rvert{{\mathbf{G}}}\lvert{\omega}\rangle\cdot{\Delta t}$ $=p_{{\Delta t}}({\omega}^{\prime}\mid{\omega})-\delta_{{\omega}}^{{\omega}^{\prime}}+o({\Delta t}),$ (A.32)
$\langle{\omega}^{\prime}_{|{\Omega}_{x}}\rvert{{\mathbf{G}}}(x)\lvert{\omega}_{|{{\mathcal{U}}}_{x}}\rangle\cdot{\Delta t}$ $=p_{{\Delta t}}({\omega}^{\prime}_{|{\Omega}_{x}}\mid{\omega}_{|{{\mathcal{U}}}_{x}})-\delta_{{\omega}_{|{\Omega}_{x}}}^{{\omega}^{\prime}_{|{\Omega}_{x}}}+o({\Delta t}).$ (A.33)
@@
$\langle{\omega}^{\prime}\rvert{{\mathbf{G}}}\lvert{\omega}\rangle\quad\text{and}\quad\langle{\omega}^{\prime}\rvert\left(\sum_{x}{{\mathbf{G}}}(x)\right)\lvert{\omega}\rangle$ (A.35)
@@
$\langle{\omega}^{\prime}\rvert{{\mathbf{G}}}\lvert{\omega}\rangle\cdot{\Delta t}$ $=p_{{\Delta t}}({\omega}^{\prime}\mid{\omega})-\delta_{{\omega}}^{{\omega}^{\prime}}+o({\Delta t})=\prod_{x}p_{\Delta t}({\omega}^{\prime}_{|{\Omega}_{x}}\mid{\omega}_{|{{\mathcal{U}}}_{x}})-\delta_{{\omega}}^{{\omega}^{\prime}}+o({\Delta t}).$ (A.36)
@@
$=\prod_{x}\left(\langle{\omega}^{\prime}_{|{\Omega}_{x}}\rvert{{\mathbf{G}}}(x)\lvert{\omega}_{|{{\mathcal{U}}}_{x}}\rangle\cdot{\Delta t}+\delta_{{\omega}_{|{\Omega}_{x}}}^{{\omega}^{\prime}_{|{\Omega}_{x}}}\right)-\delta_{{\omega}}^{{\omega}^{\prime}}+o({\Delta t})$ (A.38)
$=\sum_{x}\langle{\omega}^{\prime}_{|{\Omega}_{x}}\rvert{{\mathbf{G}}}(x)\lvert{\omega}_{|{{\mathcal{U}}}_{x}}\rangle\cdot{\Delta t}\cdot\delta_{{\omega}_{|{\Omega}/{\Omega}_{x}}}^{{\omega}^{\prime}_{|{\Omega}/{\Omega}_{x}}}+o({\Delta t}).$ (A.39)
@@
$=\sum_{x}\left(\langle{\omega}^{\prime}_{|{\Omega}_{x}}\rvert\otimes\langle{\omega}^{\prime}_{|{\Omega}/{\Omega}_{x}}\rvert\right)({{\mathbf{G}}}(x)\lvert{\omega}_{|{{\mathcal{U}}}_{x}}\rangle)\otimes\lvert{\omega}_{|{\Omega}/{\Omega}_{x}}\rangle\cdot{\Delta t}+o({\Delta t})$ (A.41)
$=\sum_{x}\langle{\omega}^{\prime}\rvert{{\mathbf{G}}}(x)\lvert{\omega}\rangle\cdot{\Delta t}+o({\Delta t}).$ (A.42)
@@
$\langle{\omega}^{\prime}\rvert{{\mathbf{G}}}\lvert{\omega}\rangle\cdot{\Delta t}=\langle{\omega}^{\prime}\rvert\left(\sum_{x}{{\mathbf{G}}}(x)\right)\lvert{\omega}\rangle\cdot{\Delta t}+o({\Delta t}).$ (A.43)
@@
$\langle{\omega}^{\prime}\rvert{{\mathbf{G}}}\lvert{\omega}\rangle=\langle{\omega}^{\prime}\rvert\left(\sum_{x}G(x)\right)\lvert{\omega}\rangle.$ (A.44)
@@
where the commutator ${{\mathbf{G}}}(x),{{\mathbf{G}}}(x^{\prime})]={{\mathbf{G}}}(x){{\mathbf{G}}}(x^{\prime})-{{\mathbf{G}}}(x^{\prime}){{\mathbf{G}}}(x)$.
@@
${{\mathbf{G}}}(x^{\prime}){{\mathbf{G}}}(x)\lvert{\omega}\rangle$ $={{\mathbf{G}}}(x^{\prime})\left(({{\mathbf{G}}}(x)\lvert{\omega}_{|{{\mathcal{U}}}_{x}}\rangle)\otimes\lvert{\omega}_{|{\Omega}/{\Omega}_{x}}\rangle\right)$ (A.47)
$=({{\mathbf{G}}}(x)\lvert{\omega}_{|{{\mathcal{U}}}_{x}}\rangle)\otimes({{\mathbf{G}}}(x^{\prime})\lvert{\omega}_{|{{\mathcal{U}}}_{x^{\prime}}}\rangle)\otimes\lvert{\omega}_{|{\Omega}/({\Omega}_{x}\times{\Omega}_{x^{\prime}})}\rangle$ (A.48)
$={{\mathbf{G}}}(x){{\mathbf{G}}}(x^{\prime})\lvert{\omega}\rangle.$ (A.49)
@@
[…]
@@
[…]
@@
$p_{T}({\omega}^{\prime}\mid{\omega})=\langle{\omega}^{\prime}\rvert e^{{{\mathbf{G}}}T}\lvert{\omega}\rangle=\int{\mathcal{D}}{\omega}\ \exp\left\{-\sum_{x}\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ L(w_{t,x},w_{t,x}^{+})\right\},$ (A.54)
@@
$p_{t}({\omega}^{\prime}\mid{\omega})$ $=\sum_{k=0}^{N}\sum_{{\omega}_{0:T,{\Delta t},k}}^{{\omega}\to{\omega}^{\prime}}\prod_{n=1}^{N}\langle{\omega}_{t_{n}}\rvert e^{{{\mathbf{G}}}{\Delta t}}\lvert{\omega}_{t_{n-1}}\rangle$ (A.55)
$=\sum_{k=0}^{N}\sum_{{\omega}_{0:T,{\Delta t},k}}^{{\omega}\to{\omega}^{\prime}}\cdot({\Delta t})^{k}\cdot\frac{1}{({\Delta t})^{k}}\prod_{n=1}^{N}\langle{\omega}_{t_{n}}\rvert e^{{{\mathbf{G}}}{\Delta t}}\lvert{\omega}_{t_{n-1}}\rangle$ (A.56)
$=\sum_{k=0}^{N}\sum_{{\omega}_{0:T,{\Delta t},k}}^{{\omega}\to{\omega}^{\prime}}\cdot({\Delta t})^{k}\cdot\exp\left\{-k\log\left({\Delta t}\right)+\sum_{n=1}^{N}\log\ \langle{\omega}_{t_{n}}\rvert e^{{{\mathbf{G}}}{\Delta t}}\lvert{\omega}_{t_{n-1}}\rangle\right\}$ (A.57)
@@
$-k\log\left({\Delta t}\right)+\sum_{n=1}^{N}\log\ \langle{\omega}_{t_{n}}\rvert e^{{{\mathbf{G}}}{\Delta t}}\lvert{\omega}_{t_{n-1}}\rangle,$ (A.58)
@@
$e^{{{\mathbf{G}}}{\Delta t}}=1+{{\mathbf{G}}}{\Delta t}+o({\Delta t})\quad\text{and}\quad{{\mathbf{G}}}=\lvert{\omega}^{\prime}\rangle G^{{\omega}^{\prime}}_{{\omega}}\langle{\omega}\rvert$. (A.59)
@@
$\log\ \langle{\omega}_{t_{n}}\rvert e^{{{\mathbf{G}}}{\Delta t}}\lvert{\omega}_{t_{n-1}}\rangle=\log\left(1+G^{{\omega}_{t_{n}}}_{{\omega}_{t_{n-1}}}\cdot{\Delta t}+o({\Delta t})\right)=G^{{\omega}_{t_{n}}}_{{\omega}_{t_{n-1}}}\cdot{\Delta t}+o({\Delta t}).$ (A.60)
For the second case, i.e., ${\omega}_{t_{n}}\neq{\omega}_{t_{n-1}}$, there are exactly $k$ jumps for the given path. Therefore, we distribute the $k\log({\Delta t})$ into this case and obtain
$-\log({\Delta t})+\log\ \langle{\omega}_{t_{n}}\rvert e^{{{\mathbf{G}}}{\Delta t}}\lvert{\omega}_{t_{n-1}}\rangle$ $=\log\left(\langle{\omega}_{t_{n}}\rvert e^{{{\mathbf{G}}}{\Delta t}}\lvert{\omega}_{t_{n-1}}\rangle/{\Delta t}\right)$ (A.61)
@@
When taking ${\Delta t}\to 0$, or equivalently $N=T/{\Delta t}\to\infty$, the $\sum_{n=1}^{N}{\Delta t}$ above becomes an integral $\int_{0}^{T}{\mathop{}\!\mathrm{d}t}$. The last term becomes $o({\Delta t})/{\Delta t}\to 0$. Then, ${\omega}_{t}$ is discontinuous w.r.t. time but with only finite discontinuous points, i.e., jumps between configurations. Since there are only finite jumps, the first term becomes just $G^{{\omega}_{t}}_{{\omega}_{t}}$, i.e.,
@@
[…]
@@
$\lim_{{\Delta t}\to 0}-k\log\left({\Delta t}\right)+\sum_{n=1}^{N}\log\ \langle{\omega}_{t_{n}}\rvert e^{{{\mathbf{G}}}{\Delta t}}\lvert{\omega}_{t_{n-1}}\rangle=-\int_{0}^{T}{\mathop{}\!\mathrm{d}t}\ L(w_{t},w_{t}^{+}).$ (A.67)
Thus, adopting the path integral notation, we find the path integral formulation of the propagation probability
$p_{T}({\omega}^{\prime}\mid{\omega})$ $=\lim_{{\Delta t}\to 0}\sum_{k=0}^{N}\sum_{{\omega}_{0:T,{\Delta t},k}}^{{\omega}\to{\omega}^{\prime}}\cdot({\Delta t})^{k}\cdot\exp\left\{-k\log\left({\Delta t}\right)+\sum_{n=1}^{N}\log\ \langle{\omega}_{t_{n}}\rvert e^{{{\mathbf{G}}}{\Delta t}}\lvert{\omega}_{t_{n-1}}\rangle\right\}$ (A.68)